# SVInvNet: A Densely Connected Encoder-Decoder Architecture for Seismic Velocity Inversion

### v1

@@ III-B3 Decoder Network
The output of this architecture is compared with the ground truth, and the disparity between them is computed using the loss function. The loss function should be capable of capturing the disparity among the velocity values as well as the shapes of the interfaces and geological structures between two velocity models. Hence, in this section of the architecture, we employ both mean absolute error (L1) and structural similarity index measure (SSIM) as loss functions, and we assign weights to each of them using two coefficients ($\lambda_{L1}$ and $\lambda_{SSIM}$ ) to enhance the derivatives of the loss functions.
IV Experiment and Results
We generate 18,000 data samples in total, with 3,600 models for each group. For instance, in a model with seven layers there are 3 different subgroups: dense layered, faulty and salt dome; 1,200 models for each subgroup.
@@ IV Experiment and Results
Since the layer count in the dataset range from 4 to 8, the total number of models can be computed as 5$\times$3$\times$1200. From each of the 15 different types of data, 800 models are randomly selected as part of the test benchmark. Hence, the total number of test dataset size is 12,000. The remaining models in the dataset (i.e. 6,000 models) are used as the training dataset. To investigate the impacts of varying the quantity of training data on the mapping or seismic inversion process, we prepared five different training datasets. Accordingly, we selected various quantities of subgroups, as 50, 100, 200, 300, and 400 models from each subgroup. Consequently, we undertook five separate training processes with 750 (15$\times$50), 1,500, 3,000, 4,500, and 6,000 training dataset sizes and evaluated the trained deep models using the same test benchmark that we prepared using 12,000 models. For the purposes of clarity and convenience, we introduce a nomenclature to denote the various training datasets. Specifically, we assign Roman numerals I-V to represent training datasets corresponding to sample sizes of 750, 1500, 3000, 4500, and 6000, respectively. Hence, these datasets will be referred to as TD-I, TD-II, TD-III, TD-IV, and TD-V, respectively.
@@
The number of test data is at least twice that of the training datasets, enabling an investigation into the generalization ability of the trained deep model. For example, TD-IV contains 300 models of each velocity subgroups, in terms of number of samples, it corresponds to 37.5% of the test benchmark size. The training conditions remained identical across all five training processes. Employing a consistent nomenclature for the trained models, we extend the use of Roman numerals I-V to the designation of our train models. For instance Model-I will be corresponding to the model that is trained with TD-I and so on.
TABLE I: Training Hyperparameters
Hyperparameter Value
Optimizer Adam
Initial learning rate 5.0 e-3
Epoch 500
Batch size 32
@@
For every velocity model, 20 shot gathers with 34 receivers are computed for a 1-second duration with a 1ms temporal sampling interval. The source interval is 35 m and the receiver interval is 21 m. Consequently, the size of the input data is [20, 1000, 34], while the size of the ground truth data (velocity model) is [1, 100, 100]. We established the network hyperparameters based on the results of various experiments. For optimization, the Adam optimizer [35] is utilized with a batch size of 32, and the initial learning rate is set to 5e-3. During the training, the learning rate is adjusted by gradually decreasing it in accordance with the optimization process. The number of epochs for all training processes is set to 500. The training hyper parameters of the deep model are depicted in Table I.
@@
Figure 6: Proposed SVInvNet architecture for seismic velocity inversion.
TABLE II: The Performances of the Three Versions of InversionNet on the Test Dataset, trained using Noiseless TD-V.
InversionNet($\alpha$) InversionNet($\beta$) InversionNet($\gamma$)
Metric L1 L1 + SSIM L1 L1 + SSIM L1 L1 + SSIM
L1 0.008797 0.008828 0.008574 0.008385 0.007249 0.007316
L2 0.000497 0.000499 0.000485 0.000471 0.000407 0.000401
SSIM 0.999970 0.999970 0.999972 0.999973 0.999977 0.999978
MSSIM 0.999997 0.999997 0.999997 0.999997 0.999997 0.999997
@@ IV-A Baseline Model
To compare the performance of SVInvNet, we employed the InversionNet model, proposed by Wu et al. [11], as the baseline. Li et al.[12] also used the InversionNet as the baseline, with certain modifications made to its architecture. Despite slight disparity, the architecture introduced by Li et al. [12] yields better results than Wu et al. version in our experiments. The parameter count for both versions of InversionNet is around 44 million.
We also made certain modifications to the InversionNet, resulting in a new version that outperforms the previous iterations (Table II). For convenience, we use $\alpha$, $\beta$, and $\gamma$ to denote Wu et al. [11], Li et al. [12], and our variant of InversionNet models, respectively. The tabulated data in Table III shows the parameter counts for each of the three iterations of both the baseline and SVInvNet.
As for the modifications, we replaced the MaxPooling function with a Convolution function in the fourth layer of the InversionNet($\alpha$) and decreased the feature size until it reached 4$\times$4 instead of 1$\times$1. Furthermore, by substituting each deconvolution layer with an upsample + convolution approach in the decoder segment of the model, the parameter count decreased to 20 million, leading to an enhancement in the prediction capabilities of the new version.
TABLE III: Model Sizes
Deep Model Parameter Count(M)
InversionNet($\alpha$) 44
InversionNet($\beta$) 44
InversionNet($\gamma$) 20
SVInvNet 4
@@ IV-B Results and Discussion
The three variants of InversionNet are trained with the TD-V separately and tested with our test benchmark. In order to examine the implications of the loss function, the training stage is conducted twice, once incorporating the L1 loss function and once integrating both the L1 and (1-SSIM) loss function depicted as L1+SSIM. The statistics of misfit and similarity metrics in the test stage are presented in the Table II.
As per the findings depicted in Table II, InversionNet($\gamma$) demonstrates superior capability in acquiring the mapping of seismic data to the velocity model in quantitative comparison to versions $\alpha$ and $\beta$. Additionally, the L1+SSIM loss function effectively captures disparity between the output of the deep model and the ground truth, surpassing the performance of the sole L1 loss function in this aspect.
Figure 7: The L1 (Left) and SSIM (Right) loss curves on the training and validation datasets obtained during the training of Model-I-III-V.
Additional experiments are carried out on the noisy dataset utilizing the L1+SSIM loss function. In this context, we utilized the noisy TD-V to train three versions of the InversionNet separately. Subsequently, we employed the noisy test benchmark to establish the testing process. The performances of the models using four metrics for the noisy test set are presented in Table IV. As can be seen, our improved version of InversionNet demonstrates better performance over InversionNet($\alpha$) and similar performance to InversionNet($\beta$). Consequently, we utilized the InversionNet($\gamma$) as the baseline for our benchmark model evaluations for the upcoming experiments.
TABLE IV: The Performances of the Three Versions of InversionNet on the Test Dataset, trained using Noisy TD-V.
InversionNet ($\alpha$) ($\beta$) ($\gamma$)
L1 0.014378 0.013028 0.013142
L2 0.000765 0.000706 0.000713
SSIM 0.999943 0.999950 0.999949
MSSIM 0.999994 0.999995 0.999995
@@
Utilizing the noiseless and noisy versions of our TD-I to TD-V, we independently trained SVInvNet and subsequently evaluated each of the resultant five trained deep models (Model-I to Model-V) using our test benchmark, which is identical for all models with 12,000 data pairs. The misfit and similarity values of the test phases for the noiseless dataset are shown in the Table V. As anticipated, increasing the volume of the trained dataset corresponded to a decrease in the values of L1 and L2, alongside an increase in the values of SSIM and MSSIM. The L1 and SSIM loss function curves on the training and validation datasets during the training are depicted for three selected models aggregated in Fig. 7. From the plots, it is evident that Model-I exhibits overfitting behavior. However, as the training dataset size increases (refer to Model-V), this issue is significantly reduced.
@@
Increasing the number of samples in the training dataset resulted in a decrease in loss function values, an improvement in similarity measures, and an extension of the training phase duration. Based on the results presented in Tables II and V, it is evident that the performance of our proposed deep model surpasses that of the InversionNet. Even the loss values of SVInvNet, trained with TD-III, indicate a superior performance compared to the values of the InversionNet($\gamma$), trained with TD-V. In other words, our deep model can map seismic shot gathers to velocity models better with lower loss function than the baseline model.
Qualitative evaluations of our trained models are depicted for randomly selected test samples in Fig. 8. Each sample is purposefully selected from 8-layer group to showcase the visuals from more challenging group with respect to the inversion for layered, fault, and salt dome categories. The ground truth and the predicted result of Model-I, Model-III, and Model-V of these samples are depicted in Fig. 8. For qualitative assessments, we pay attention to the velocity value and the interface shape of the layers, the fault line continuity, and throw value as well as the shape, position, and velocity value of the salt dome. As can be seen from the samples, the velocity values within the layers remain constant. With an increase in the size of the training dataset, the discrepancies in the approximation of this value diminish, gradually converging towards a stable and unvarying value. Evident from the velocity profile depicted in the same figure on the right column, the blue line, associated to Model-I, exhibits more pronounced oscillations compared to the green line (corresponding to Model-III), while the green line displays greater oscillatory behavior than the red line (Model-V). Notably, minimal oscillation is observable in the red line.
Figure 8: Ground truth velocity model and the prediction results of our Model-I (the second column), Model-III (the third column), and Model-V (the fourth column). The rightmost column shows a vertical velocity profile at the gray line depicted in the ground truth model for four velocity models.
@@
TABLE V: The Performance of Our Deep Models on the Test Set. Models I-V are trained using Noiseless TD I-V, respectively.
Model-I Model-II Model-III Model-IV Model-V
L1 0.013680 0.009833 0.006844 0.005301 0.004944
L2 0.000737 0.000509 0.000356 0.000272 0.000253
SSIM 0.9999436 0.9999674 0.9999810 0.9999870 0.9999884
MSSIM 0.99999491 0.99999692 0.99999812 0.99999870 0.99999882
@@
Figure 9: Ground truth velocity model and prediction results of InversionNet $\beta$ and $\gamma$ version and our Model-V. The rightmost column shows a vertical velocity profile at the gray line depicted in the ground truth model for four velocity models. The white dashed lines in the models of the first row are employed to illustrate the local position of the second and third interfaces. The yellow arrow and white rectangular within the faulty velocity model of Model-V are employed to highlight the alteration of the third interface and the estimation of impact of the fault on the sixth interface, respectively.
TABLE VI: The Performance of Our Deep Models on the Test Set. Models I-V are trained using Noisy TD I-V, respectively.
Model-I Model-II Model-III Model-IV Model-V
L1 0.024961 0.018942 0.014189 0.010378 0.008034
L2 0.001548 0.001044 0.000736 0.000489 0.000399
SSIM 0.999856 0.999911 0.999945 0.999967 0.999977
MSSIM 0.999988 0.999992 0.999995 0.999997 0.999997
For visual comparison of the performance between InversionNet $\beta$ and $\gamma$ version and our Model-V, three samples selected randomly from seven layered models are illustrated in Fig. 9. The deep models are trained with noiseless TD-V. In the layered velocity model depicted in the first row of Fig. 9, the second and third interfaces display curvature, and the velocity contrast between the third and fourth layers is relatively minimal. As a result, the amplitude and trajectory of the reflected hyperbolas from the second and third interfaces are weakened and demonstrate greater complexity compared to the other interfaces. The white dashed line delineates the locations of the second and third interfaces at a specific point. The results from our developed model surpass those of InversionNet($\gamma$), and InversionNet($\gamma$)’s results, in turn, outperform those of InversionNet($\beta$). This superior performance is also maintained in samples featuring fault or salt dome.
Figure 10: Prediction outcomes of our models for the noisy dataset. The first column corresponds to the ground truth (target) velocity model, while the rightmost column displays a vertical velocity profile along the gray line depicted in the target model for four velocity models. The velocity models are the same as Fig. 8.
@@
The introduction of noise to the data exacerbates this situation, causing distortion in the amplitude of the signal emanating from the interfaces and deforming the trajectory of the scattering or reflection hyperbola. This further complicates the process of extracting features from the seismic gathers performed by CNN kernels. The adverse impact of noise in the velocity mapping process is evident based on the performances presented in Tables VI and V. Upon examining the results in Table IV and VI, it is evident that SVInvNet demonstrates superior performance, even in the presence of noise, when compared to the InversionNet different versions.
Figure 11: Comparison of FWI and Model-V results. The columns corresponds to the ground truth, initial, FWI result, and Model-V estimation velocity model respectively, while the rightmost column displays a vertical velocity profile along the gray line depicted in the ground truth model for four velocity models.
@@
Fig. 10, depicts the predicted velocity values of SVInvNet trained with noisy TD-I, TD-III, and TD-V. The performances of Model-I and Model-III show significantly inferior results compared to the corresponding noiseless cases. The prediction results presented in Table VI substantiate this, especially when compared with similar items in Table V. To successfully construct the velocity model by capturing the signal within noisy data and discerning its pattern, the deep model requires a greater number of samples for supervision. In other words, an increased sample size is imperative for the model to effectively filter out noise and accurately learn the desired signal.
IV-B1 Comparison with FWI Method
In this section, a comparison was conducted between the outcomes of the Model-V and the derivative-based FWI method using examples of velocity models featuring fault and salt dome. Fig. 11 illustrates that the velocity models derived by Model-V exhibit a closer alignment with the ground truth when contrasted with the FWI method.
Due to the high sensitivity of the FWI method to noise, noise-free datasets are used. Initial velocity model is acquired through Gaussian smoothing with a standard deviation of $5$ applied to the ground truth model. To minimize the cyclic skipping effect [9], we executed the process with the selection of five cutoff frequencies within [10, 15, 20, 25, 30] Hz. Furthermore, a condition was introduced by limiting the velocity parameters between 1000 and 5000 m/s to enhance the efficacy of the FWI method. The entire process is performed for a total of 50 iterations. The computational configuration employed for seismic shot gathers in the trained data segment remains consistent with the setup utilized in the FWI procedure. To implement FWI, the DeepWave[36] package is employed.
As illustrated in Fig. 11, FWI encounters challenges in estimating the shape of deeper layer interfaces, despite having a strong initial model. Furthermore, the velocity models between layers in the FWI estimation exhibit non-constant behavior and display oscillatory patterns. The profile section, green line representing FWI results, reveals heightened instability compared to the red line representing Model-V result.
Concerning the salt dome, while the initial velocity model provides a somewhat clear indication of the salt dome’s location and distribution, Model-V demonstrates superior performance in inverting both the shape and velocity values of the salt mass in compare with FWI method.

### v2

@@ III-B3 Decoder Network
The output of this architecture is compared with the ground truth, and the discrepancy between them is computed using the loss function. The loss function should be capable of capturing the disparity among the velocity values as well as the shapes of the interfaces and geological structures between two velocity models. Hence, in this section of the architecture, we employ both mean absolute error (L1) and structural similarity index measure (SSIM) as loss functions, and we assign weights to each of them using two coefficients ($\lambda_{L1}$ and $\lambda_{SSIM}$ ) to enhance the derivatives of the loss functions.
IV Experiments and Results
@@
Given that the number of layers in each velocity model ranges from 4 to 8, the total number of models, including layered, faulted, and salt dome types, is calculated as 5$\times$3$\times$1200. From each of the 15 different subgroups of data, 800 models are randomly selected as part of the test benchmark. Hence, the total number of test dataset size is 12,000. The remaining models in the dataset (i.e. 6,000 models) are used as the training dataset. To investigate the impacts of varying the quantity of training data on the mapping or seismic inversion process, we prepared five different training datasets. Accordingly, we selected various quantities of subgroups, as 50, 100, 200, 300, and 400 models from each type. Consequently, we undertook five separate training processes with 750 (15$\times$50), 1,500, 3,000, 4,500, and 6,000 training dataset sizes and evaluated the trained deep models using the same test benchmark that we prepared using 12,000 models. For the purposes of clarity and convenience, we introduce a nomenclature to denote the various training datasets. Specifically, we assign Roman numerals I-V to represent training datasets corresponding to sample sizes respectively. Hence, these datasets will be referred to as TD-I, TD-II, TD-III, TD-IV, and TD-V, respectively. The training datasets and related information for each are provided in Table I.
@@
The number of test data is at least twice that of the training datasets, enabling a more accurate investigation into the generalization ability of the trained deep models[40]. For example, TD-IV contains 300 models of each velocity subgroups, in terms of number of samples, it corresponds to 37.5% of the test benchmark size. The training conditions remained identical across all five training processes. Employing a consistent nomenclature for the trained models, we extend the use of Roman numerals I-V to the designation of our train models. For instance Model-I will be corresponding to the model that is trained with TD-I and so on.
@@
We established the network hyperparameters based on the results of various experiments. For optimization, the Adam optimizer [41] is utilized with a batch size of 32, and the initial learning rate is set to 5e-3. During the training, the learning rate is adjusted by gradually decreasing it in accordance with the optimization process. The number of epochs for all training processes is set to 500.
@@
TABLE I: The training datasets and their respective subgroups
Training dataset TD-I TD-II TD-III TD-IV TD-V
Sample size 750 1500 3000 4500 6000
Layer counts Ranges uniformly from 4 to 8 layers
Structural features Dense layered - Faulty - Salt dom
Seismic data 34 receivers, 20 sources, and 1s records
Noise content two datasets: noisy and noiseless
@@ IV-A Baseline Model
To compare the performance of SVInvNet, we employed the InversionNet model, proposed by Wu et al. [13], as the baseline, with certain modifications made to its architecture. As for the modifications, we replaced the MaxPooling function with a Convolution function in the fourth layer of the InversionNet and decreased the feature size until it reached 4$\times$4 instead of 1$\times$1. Furthermore, by substituting each deconvolution layer with an upsample + convolution approach in the decoder segment of the model, the parameter count decreased to 20 million, leading to an enhancement in the prediction capabilities.
@@ IV-B Results and Discussion
Utilizing the noiseless and noisy versions of our TD-I to TD-V, we independently trained SVInvNet and subsequently evaluated each of the resultant five trained deep models (Model-I to Model-V) using our test benchmark. The misfit and similarity values of the test phases for the noiseless and noisy dataset are shown in the Table II. As anticipated, increasing the volume of the trained dataset corresponded to a decrease in the values of L1 and L2, alongside an increase in the values of SSIM and MSSIM.
The L1 and SSIM loss function curves on the training and validation datasets during the training are depicted for three selected models aggregated in Fig. 7. The loss curves of the other models are bounded by Model-I and Model-V curves. From the plots, it is evident that Model-I exhibits overfitting behavior. However, as the training dataset size increases (refer to Model-V), this issue is significantly reduced.
Figure 7: The L1 (Left) and SSIM (Right) loss curves on the training and validation datasets obtained during the training of Model-I-III-V. Best viewed in the electronic version with zoom.
InversionNet is trained using both the noiseless and noisy TD-V datasets under conditions equivalent to those of SVInvNet. The statistics of misfit and similarity metrics in the test stage are presented in the Table II as InversionNet-V.
@@
Based on the results presented in Table II it is evident that the performance of our proposed deep model surpasses that of the InversionNet. Even the loss values of SVInvNet, trained with TD-III, indicate a superior performance compared to the values of the InversionNet, trained with TD-V.
@@
TABLE II: The Performance of SVInvNet on the Test Set. Models I-V are trained utilizing Noiseless and Noisy TD I-V, respectively. The Last Column Represents Results of InversionNet Trained Using Noiseless and Noisy TD-V.
Noiseless Dataset Noisy Dataset
L1 L2 SSIM MSSIM L1 L2 SSIM MSSIM
Model-I 0.013680 0.000737 0.9999436 0.99999491 0.024961 0.001548 0.999856 0.999988
Model-II 0.009833 0.000509 0.9999674 0.99999692 0.018942 0.001044 0.999911 0.999992
Model-III 0.006844 0.000356 0.9999810 0.99999812 0.014189 0.000736 0.999945 0.999995
Model-IV 0.005301 0.000272 0.9999870 0.99999870 0.010378 0.000489 0.999967 0.999997
Model-V 0.004944 0.000253 0.9999884 0.99999882 0.008034 0.000399 0.999977 0.999997
InversionNet-V 0.007316 0.000401 0.999978 0.999997 0.013142 0.000713 0.999949 0.999995
Qualitative evaluations of our trained models are depicted for randomly selected test samples in Fig. 8. Each sample is purposefully selected from 8-layer group to showcase the visuals from more challenging group with respect to the inversion for layered, fault, and salt dome categories. The ground truth and the predicted result of Model-I, Model-III, and Model-V of these samples are depicted in Fig. 8. For qualitative assessments, we pay attention to the velocity value and the interface shape of the layers, the fault line continuity, and throw value as well as the shape, position, and velocity value of the salt dome. As can be seen from the samples, the velocity values within the layers remain constant. With an increase in the size of the training dataset, the discrepancies in the approximation of this value diminish, gradually converging towards a stable and unvarying value.
Figure 8: Ground truth velocity model and the prediction results of our Model-I (the second column), Model-III (the third column), and Model-V (the fourth column). The rightmost column shows a vertical velocity profile at the gray line depicted in the ground truth model for four velocity models. Best viewed in the electronic version with zoom.
Figure 9: Prediction outcomes of our models for the noisy dataset. The first column corresponds to the ground truth (target) velocity model, while the rightmost column displays a vertical velocity profile along the gray line depicted in the target model for four velocity models. The velocity models are the same as Fig. 8. Best viewed in the electronic version with zoom.
@@
Figure 10: Sample ground truth velocity models and their prediction results of InversionNet, VelocityGAN, and SVInvNet for a) CurveVel-B and b) CurveFault-A benchmarks, with the vertical velocity profiles on the right-most column. The velocity profiles belong to the gray lines depicted in the corresponding ground truth models. Best viewed in the electronic version with zoom.
TABLE III: The Performance of Four DNN Models on the Test Set. Models are trained utilizing CurveVel-B (CV-B) and CurveFault-A (CF-A) datasets.
CV-B CF-A
Metric MAE RMSE SSIM MAE RMSE SSIM
SVInvNet 0.0250 0.0583 0.9997 0.0042 0.0128 0.9999
InversionNet 0.1448 0.3111 0.6630 0.0303 0.0766 0.9448
VelocityGAN 0.1268 0.2618 0.7111 0.0216 0.0505 0.9687
UPFWI 0.1777 0.3179 0.6614 0.0500 0.0966 0.9495
@@
The introduction of noise to the data exacerbates this situation, causing distortion in the amplitude of the signal emanating from the interfaces and deforming the trajectory of the scattering or reflection hyperbola. This further complicates the process of extracting features from the seismic gathers performed by CNN kernels. The adverse impact of noise in the velocity mapping process is evident based on the performances presented in Tables II. In the Table II, it is evident that SVInvNet demonstrates superior performance, even in the presence of noise, when compared to the InversionNet.
@@
Fig. 9, depicts the predicted velocity values of SVInvNet trained with noisy TD-I, TD-III, and TD-V. The performances of Model-I and Model-III show significantly inferior results compared to the corresponding noiseless cases. The prediction results presented in Table II substantiate this. To successfully construct the velocity model by capturing the signal within noisy data and discerning its pattern, the deep model requires a greater number of samples for supervision. In other words, an increased sample size is imperative for the model to effectively filter out noise and accurately learn the desired signal.
IV-B1 Analysis Using the OpenFWI Dataset
The performance evaluation of the SVInvNet deep model is conducted using the open source dataset OpenFWI, as published by Deng et al. [18]. They trained three distinct deep models utilizing various subgroups of OpenFWI and report the results. These deep models include InversionNet [13], VelocityGAN [42], and UPFWI [43]. We employed the subset benchmarks of OpenFWI, namely the CurveVel-B and CurveFault-A datasets, for both the training and testing phases of the SVInvNet.
Compared to our dataset, the OpenFWI dataset includes 70 receivers and uses velocity models sized at 70$\times$70. To match this, we add an extra layer to the SVInvNet first four layers and change the final layers’ size from 100$\times$100 to 70$\times$70. It was noted to emphasize that both training and testing were conducted on the same benchmark, as previously discussed. Therefore, it was removed. We compare our results with the best values from their study, as shown in Table III. The data in Table III clearly shows that our model performs better than the other deep models. We also utilized the corresponding pretrained models of InversionNet and VelocityGAN to provide visual comparisons with SVInvNet, using samples randomly selected from both the CurveVel-B and CurveFault-A test datasets, as presented in Fig. 10. The samples in Fig. 10 align with the quantitative values presented in Table III. SVInvNet demonstrated superior performance in predicting both the velocity values and the interface shapes when compared to the other models.
Figure 11: Sample ground truth velocity models and the corresponding prediction results from the baseline model and SVInvNet for the Marmousi benchmarks, with vertical velocity profiles shown in the right-most column. The velocity profiles correspond to the violet, gray, and yellow lines as shown in their respective ground truth models.
IV-B2 Analysis Using the Marmousi-based models
To assess the performance of the SVInvNet, the models obtained from the Marmousi model was utilized, as it provides a more accurate representation of the true velocity distribution within the subsurface. To accomplish this, we generated 407 sub-model sets by systematically scanning the Marmousi velocity model with a 200×200 template at predetermined intervals. We ensured that the velocity contrast in the prepared models was at least 200 m/s in the vertical direction. The models were then downsampled to a 100×100 size. We randomly selected 23 models from the dataset for testing, while the remaining 384 models were combined with our own TD-V to train the baseline model and SVInvNet. The trained models was then tested on the 23 test datasets. Qualitative evaluations of the results for randomly selected test samples are presented in Fig. 11. As demonstrated in the Fig. 11, SVInvNet outperforms the baseline; however, the training dataset is inadequate for accurately estimating high-slope faulted transitions and nested sloped layers with low contrast.